OpenStudy (anonymous):

Simplify. q ^ 33/4 over q ^8

OpenStudy (anonymous):

\[\frac{q^{\frac{33}{4}}}{q^{8}} = q^{\frac{33}{4} - 8}\]

OpenStudy (anonymous):

q^1/4

OpenStudy (anonymous):

thank u!

OpenStudy (anonymous):

np :)